# The representation of Convectively Coupled Equatorial Waves and upscale energy transfer in models with explicit and parametrized convection

### v1

@@ 5 Conclusions
The general conclusions are that the RAL (convection permitting) and GAL (convection parametrized) models simulate similar dynamical structures of equatorial waves, albeit with differences in vertical structures and vertical propagation. RAL simulates 56% more upscale energy transfer than GAL in the troposphere, with similar magnitudes of downscale transfer. We hypothesis that these large differences in energy transfer are because RAL explicitly represents horizontal motions associated with deep convection, whereas the parametrization scheme in GAL acts to remove vertical instability in a column without the associated convective circulation.
@@
RAL also exhibits higher correlations between this transfer and the dynamical wave, which we hypothesise is as a result of propagating convection travelling with the waves. The correlations between interscale energy transfer and the dynamical wave peak between length scales of 440 and 1760 km in the Kelvin wave and these correlations are up to 12% higher in RAL than GAL. This is accompanied by 20% higher covariance in the RAL, reflecting the higher amplitudes of upscale energy transfer. In the R1 wave, interscale energy transfer is most correlated with wave vorticity at $L=880$ km, with 10% higher correlation in RAL, and comparable covariances. In the WMRG wave, due to its large wavelength, there is no discernible peak in correlation or covariance between interscale energy transfer and wave meridional flow at the largest scales accessible in these simulations. The correlations and covariances increase with increasing scale, with correlations averaging 30% higher in RAL in the scale range 440 - 1760 km, and covariances averaging 100% higher.
@@
These results have shown that CPMs simulate more scale coupling between the small scales and large-scale dynamics in the tropical atmosphere. We hypothesise that this increased coupling is due to the explicit representation of convection. The conclusions drawn here are however limited in that only one simulation of 40 days for each model was analysed. It has been shown that RAL simulates more strongly convectively coupled EWs than GAL through the proxy of propagating interscale energy transfer and precipitation. Although the dynamical waves at any given level may appear similar in amplitude and abundance between the models (Fig. 2), the propagating convection through either of the proxies used is very different between the two models, with RAL exhibiting up to 30% more convective coupling. Rigorously demonstrating that interscale energy transfer propagating with an equatorial wave is an effective proxy for convective coupling is beyond the scope of this study, and left to future investigation. It has also been shown that the equatorial waves simulated by the two models are typically propagating too fast, and with a higher degree of variability compared to the UM analysis. This variability is larger in GAL than in RAL for CCKWs and the WMRG wave, but smaller for R1 waves.
@@
We find that interscale energy transfer propagates along with CCEWs, with the RAL model simulating more propagation and deeper, more coherent structures in interscale energy transfer. These propagating features were investigated by the superposition of wave-filtered flows over interscale energy transfer in an instantaneous (Fig. 7) and wave-following composite (Fig. 8) visualisation. It is found that in the wave-comoving frame, upscale energy transfer is associated with Kelvin wave divergence, R1 cylonic vorticity, and WMRG southward cross-equatorial meridional velocity.
@@
It has also been shown that in a time-mean sense, the RAL model simulates more upscale and less downscale than GAL over the tropics, with RAL being net upscale over the DYAMOND Summer study period in the upper troposphere, and GAL net downscale. RAL exhibits higher spatial variance in interscale energy transfer than GAL. This time mean behaviour in local interscale energy transfer is reflected in the large-scale circulation, with upscale being associated with deep convection in the ITCZ, shallow convection over rainforests, and monsoon flow. Downscale associated with the subtropical jet is seen in both models and they exhibit similar time-mean Hadley cell circulation.

### v2

@@ 5 Conclusions
The general conclusions are that the RAL3 (convection permitting) and GAL9 (convection parametrized) models simulate similar dynamical structures of equatorial waves, albeit with differences in vertical structures and vertical propagation. RAL3 simulates 56% more upscale energy transfer than GAL9 in the troposphere, with similar magnitudes of downscale transfer. We hypothesise that these differences in energy transfer are because RAL3 explicitly represents horizontal motions associated with deep convection, whereas the parametrization scheme in GAL9 acts to remove vertical instability in a column without the associated convective circulation.
@@
RAL3 also exhibits higher correlations between this transfer and the dynamical wave, which we hypothesise is as a result of propagating convection travelling more coherently with the waves (higher degree of convective coupling). The correlations between interscale energy transfer and the dynamical wave peak at $L=880$ km in the Kelvin wave, and this correlation is 8% higher in RAL3 than GAL. The ‘amplitude’ of the trough down to the maximum correlation is also more prominent in RAL3 than in GAL9 by eye. This is accompanied by 80% higher covariance in RAL3, reflecting the higher amplitudes of upscale energy transfer.
In the R1 wave, interscale energy transfer is most correlated with wave vorticity at $L=880$ km in RAL3 and at $L=440$ km in GAL9. The maximum correlation in RAL3 for the R1 wave is 48% higher than in GAL9. This pattern is repeated in the covariances, with RAL3 maximum covariance between the wave and R1-following interscale energy transfer being 60% higher than in GAL9.
In the WMRG wave, due to its large wavelength, there is no discernible peak in correlation or covariance between interscale energy transfer and wave meridional flow at the largest scales accessible in these simulations. The correlations and covariances increase with increasing scale, with correlations averaging 85% higher in RAL3 across all scales, and covariances averaging 171% higher.
@@
These results have shown that CPMs simulate more scale coupling between the small scales and large-scale dynamics in the tropical atmosphere. We hypothesise that this increased coupling is due to the explicit representation of convection. The conclusions drawn here are however limited in that only one simulation of 40 days for each model was analysed. It has been shown that RAL3 simulates more strongly convectively coupled EWs than GAL9 through the proxy of propagating interscale energy transfer and precipitation. Although the dynamical waves at any given level may appear similar in amplitude and abundance between the models (Fig. 3), the propagating convection through either of the proxies used is very different between the two models, with RAL3 exhibiting an average of 47% more convective coupling. Rigorously demonstrating that interscale energy transfer propagating with an equatorial wave is an effective proxy for convective coupling is beyond the scope of this study, and left to future investigation. It has also been shown that the equatorial waves simulated by the two models are typically propagating too fast, and with a higher degree of variability compared to the UM analysis. This variability is larger in GAL9 than in RAL3 for CCKWs and the WMRG wave, but smaller for R1 waves.
@@
We find that interscale energy transfer propagates along with CCEWs, with the RAL3 model simulating more propagation and deeper, more coherent structures in interscale energy transfer. These propagating features were investigated by the superposition of wave-filtered flows over interscale energy transfer in an instantaneous (Fig. 8) and wave-following composite (Fig. 9) visualisation. It is found that in the wave-comoving frame, upscale energy transfer is associated with Kelvin wave divergence, R1 cyclonic vorticity, and WMRG southward cross-equatorial meridional velocity.
@@
It has also been shown that in a time-mean sense, the RAL3 model simulates more upscale and less downscale than GAL9 over the tropics, with RAL3 being net upscale over the DYAMOND Summer study period in the upper troposphere, and GAL9 net downscale. RAL3 exhibits higher spatial variance in interscale energy transfer than GAL. This time mean behaviour in local interscale energy transfer is reflected in the large-scale circulation, with upscale being associated with deep convection in the ITCZ, shallow convection over rainforests, and monsoon flow. Downscale associated with the subtropical jet is seen in both models and they exhibit similar time-mean Hadley cell circulation.